OpenStudy (anonymous):

does any one know how they could tell me how to write a problem when you have a number with smaller number which would make it and number square such as 3x3=

OpenStudy (anonymous):

Not sure what your asking here. Your grammar is a little sparse. Give an example question. 3x3=9 haha, but I am sure that is not what you are looking for....

OpenStudy (anonymous):

I'm going to try here: what I am trying to say is 3x3=6 because what you wrote would be this 3x3x3= 9 in what I am trying to explain.

OpenStudy (anonymous):

Sorry I dont know what your asking. \[3*3 = 9\ne 6 \] are you looking for numbers that add up and multiply to form the same number? like \[2x2 = 2+2 = 4\]